OpenStudy (anonymous):

Find the hypotenuse of each isosceles right triangle when the legs are of the given measure. Given =

Nnesha (nnesha):

a^2+b^2=c^2 apply the Pythagorean theorem

OpenStudy (anonymous):

im not soo good in math, but i really want to undestand how to find the answer

Nnesha (nnesha):

given =what ?

OpenStudy (anonymous):

6 sqare root of 2

OpenStudy (anonymous):

i have to find the inches

OpenStudy (anonymous):

would it be 8.49? im really just guessing, someone help please????

OpenStudy (anonymous):

An isosceles right triangle is one where the legs are the same. The hypotenuse is √2 times the length of the legs in this type of triangle. So multiply your given length by √2

OpenStudy (anonymous):

i get 8.48528137424 do i have to keep going? or am i doing this all wrong? @peachpi

OpenStudy (anonymous):

I think they want an exact answer. 8.49 is just an approximation for the given 6√2. Multiply 6√2 by √2

OpenStudy (anonymous):

36

OpenStudy (anonymous):

\[6\sqrt{2} \times \sqrt{2}=6\sqrt{2\times2}=?\]

OpenStudy (anonymous):

it is 12 sorry

Nnesha (nnesha):

or you can use the equation \[\huge\rm (6\sqrt{2})^2+(6\sqrt{2})^2=c^2\]

OpenStudy (anonymous):

yes 12

OpenStudy (anonymous):

jeje sorry i was confused but i understand now thank you soo much for your help @Nnesha and @peachpi!!!!

OpenStudy (anonymous):

you're welcome

Nnesha (nnesha):

np :=)